refer to the following information.
About 60% of adults get less than the recommended 8 hours of sleep per night according to a study conducted by the University of Washington.
In a random sample of 7 adults,
a) what is the expected number of them that get less than the recommended 8 hours of sleep per night? (That is, find the mean.)
b) what is the standard deviation of this distribution?

Answer by stanbon(57362)

About 60% of adults get less than the recommended 8 hours of sleep per night according to a study conducted by the University of Washington.
In a random sample of 7 adults,
a) what is the expected number of them that get less than the recommended 8 hours of sleep per night? (That is, find the mean.)
mean = np = 7*0.6 = 4.2
--------
b) what is the standard deviation of this distribution?
std = sqrt(npq) = sqrt(4.2*0.4) = 1.2961
====================
Cheers,
Stan H.
=============